In a random sample of 84 coffee shops in Vancouver, it is found that a latte costs an average of $4.65 with a standard deviation of $0.30. Using alpha = 0.05, determine if the average price is significantly higher than $4.50.

1.Fill in the blanks to construct the correct null and alternate hypotheses.

H0:                            [ Select ]                       ["p", "p-hat", "μ", "y-bar"]                                    [ Select ]                       ["<", "=", ">", "≠"]                                    [ Select ]                       ["4.50", "4.65"]      

HA:                            [ Select ]                       ["p", "p-hat", "μ", "y-bar"]                                    [ Select ]                       ["<", "=", ">", "≠"]                                    [ Select ]                       ["4.50", "4.65"]      

2.What is the value of the test statistic?

3.If a p-value of 0.0054 were found, what would the decision be? Fill in the blanks to make the sentence below correct.

                           [ Select ]                       ["Fail to reject H0.", "Reject H0."]         The average is                            [ Select ]                       ["significantly different from", "significantly less than", "significantly more than"]                                    [ Select ]                       ["$4.50", "$4.65"]         .
Answer #1

1) H0 : µ = 4.5

Ha : µ > 4.5

2) Test statistic =\frac{\bar{X}-\mu}{s/\sqrt{n}}

                         =\frac{4.65-4.5}{0.3/\sqrt{84}}

                         = 4.58

3) Since p-value is less than alpha, we should reject the null hypothesis.

Reject H0. The average is significantly more than $4.50